is 9.1 M HCL and electrolyte


1
Expert's answer
2022-04-18T17:03:03-0400

Electrolytes are substances, melts, or solutions in which an electric current is generated by the motion of ions rather than electrons, as in ordinary conductors (although it is possible to combine both types of conductivity in one substance). Ions are formed due to electrolytic dissociation - the decomposition of complex substances into ions.

C(H+) = C(HCl) = 9.1 M;

pH = -log(C(H+)) = -log(9.1) = -0.96.